Determine the equation of the parabola with vertex (-2,5) and that passes through (4,-8)
Answer by stanbon(75887)

Determine the equation of the parabola with vertex (-2,5) and that passes through (4,-8)
---------------
Vertex form of a parabola:
(4p)(y-k) = (x-h)^2
Substitute to find "4p"
4p(-8-5) = (4+2)^2
4p(-13) = 36
4p = -36/13
--------------
EQUATION:
(-36/13)(y-5)=(x+2)^2
Multiply both sides by 13 to get:
-36(y-5) = 13[x^2+4x+4]
-36y + 180 = 13x^2+52x+52
-36y = 13x^2+52x-128
y = -13/36)x^2-(13/9)x+(13/9)
===============================
Cheers,
Stan H.